Please help me solve this equation: And I need to know how to check the answer, so ...solve and check
3x²-2x+7-x²+3-2x²=4
Thanks, in advance

Answer by rfer(16322)

-2x=4-10
-2x=-6
x=-6/-2
x=3
-----------------
3x^2-2*3+7-x^2+3-2x^2=4
-6=-6